# Deformation and breakup of the liquid ligament with various disturbances on the interface in shear flow

### v1

@@ Abstract
This study performed a numerical analysis of the deformation and breakup of a liquid ligament with various disturbances on the interface in shear flow. The shear flow generates a three-dimensional flow and vortices around the liquid ligament. These vortices promote the movement of the liquid inside the liquid ligament. When the velocity difference of shear flow increases, a nonlinear effect becomes strong, and turbulence with higher wavenumber components than the initial disturbance occurs at the interface. This turbulence accelerates the ligament splitting and increases the number of breakup droplets. Then, the droplet diameters become uniform, and the atomization quality improves. As the wavenumber of the disturbance applied to the interface increases, the liquid moving velocity along the central axis of the liquid ligament increases. As the wavenumber increases, the breakup time of the liquid ligament becomes short. In addition to the initial reference disturbance with a low wavenumber, when turbulence with twice the wavenumber of the reference disturbance is applied to the interface, the interface deformation and the splitting of the liquid ligament are similar to those with the single reference disturbance. When turbulence with four times the wavenumber is added to the initial disturbance with the low wavenumber at the interface, the deformation of the liquid ligament is accelerated, and the liquid ligament splits faster. In addition, the total surface area of the liquid increases because the number of breakup droplets increases. The droplet diameter becomes uniform, and therefore, the atomization quality of the liquid ligament is improved.
@@ 1 INTRODUCTION
In the general atomization process, as the ejected liquid in the form of a liquid sheet or column is destabilized by the shearing of gas and liquid, a fine-scale liquid column, called a ligament, is generated, and the liquid eventually splits into small-scale droplets (Hashimoto, 1995). Numerous studies have been reported on the destabilization of liquid films, liquid columns, and droplets (Squire, 1953; O’Rourke and Amsden, 1987; Li and Tankin, 1991; Li et al., 2000; Kourmatzis and Masri, 2014; Fernandez et al., 2009; Shinjo and Umemura, 2010; Li et al., 2011; Odier et al., 2015; Gong et al., 2015). As spray consists of many small-scale droplets, it is considered that the behavior and splitting of liquid ligaments that generate droplets significantly affect the spray characteristics. However, the destabilization and breakup of liquid ligaments have not been elucidated in detail, and therefore, further research focusing on the behavior of liquid ligaments is necessary.
@@
Marmottant and Villermaux (2004a) experimentally investigated the splitting behavior of liquid ligaments formed by pulling up a circular tube immersed in liquid and stretching the liquid. It has been clarified from this investigation that the ligament breakup time depends on the diameter of the tube, not on the speed with which the liquid is pulled up. Bizjan et al. (2014) conducted experiments on the breakup of liquid ligaments formed by centrifugal force using a rotating disk. As a result, it has been clarified that the Weber number, Ohnesorge number, and liquid volumetric flow rate greatly influence the behavior and splitting length of the liquid ligament. In addition to the above studies, experimental (Henderson et al., 2000) and numerical (Dumouchel et al., 2015; Zhang and Yuan, 2022) studies on liquid ligaments have been reported. These studies did not focus on the deformation of liquid due to gas–liquid shear velocity and did not investigate the effect of surrounding airflows on the behavior of liquid ligaments. Constante-Amores et al. (2021) captured a ligament around a liquid jet interface with direct numerical simulation. The ligament detached from the interface and formed droplets. In this manner, fine-scale liquids such as ligaments split in atomization processes. Therefore, the breakup characteristics of ligaments should be investigated to clarify spray characteristics. The authors (Yanaoka and Nakayama, 2022) performed a three-dimensional analysis of the destabilization and splitting of a liquid ligament in shear flow. The breakup time of the ligament in the shear flow became shorter as the wavenumber of initial disturbance increased, and the diameter of the formed droplet became small. As the velocity difference of the shear flow increased, the dispersion of the droplet diameter decreased, and the droplet size became uniform.
@@
Airflows exist around liquids injected from injectors. As liquid ligaments are destabilized by gas–liquid shear velocity, it is necessary to investigate the effect of shear flow on the deformation and splitting of liquid ligaments. In addition, both air and liquid flows are turbulent in high-speed jets. In such a turbulent flow field, the velocity turbulence intensity increases around the gas–liquid interface. At this time, disturbances with various wavenumber components occur at the interface, and it is considered that these turbulences promote the deformation and splitting of liquid ligaments. However, it is difficult to experimentally investigate the effects of various turbulences on interface deformation, focusing on fine-scale liquid ligaments generated in a turbulent flow. Therefore, the influences of wide-ranging wavenumber disturbances on interface deformation have not been fully elucidated. Thus, it is necessary to investigate the effects of various disturbances on the behavior of liquid ligaments.
@@ 2 NUMERICAL PROCEDURES
This study considers incompressible viscous flow. We apply the level set method (Sussman et al., 1994) to track the interface. In addition, the continuum surface force method (Brackbill et al., 1992) is used to evaluate surface tension, which is treated as a body force. The governing equations for incompressible two-phase flow are the continuity equation, the Navier-Stokes equation, the advection equation for the level set function, and the reinitialization equation, which are given as follows:
@@
where $t$ is the time, ${\bf u}=(u,v,w)$ is the velocity vector at the coordinates ${\bf x}=(x,y,z)$, $\rho$ is the density, $p$ is the pressure, $\mu$ is the viscosity coefficient, and ${\bf D}$ is the strain rate tensor. $\kappa$ is the interface curvature, $\hat{\bf n}$ is the unit normal vector at the interface, $\delta_{s}$ is the delta function, and $\phi$ is the level set function. $\tau$ is a pseudo time for reinitialization, and $\phi_{0}$ is the level set function before reinitialization. The variables in the fundamental equations are non-dimensionalized as follows using the reference values of length $L_{\mathrm{ref}}$, velocity $U_{\mathrm{ref}}$, density $\rho_{\mathrm{ref}}$, and viscosity coefficient $\mu_{\mathrm{ref}}$:
@@
The numerical method is almost the same as the previous research (Yanaoka and Nakayama, 2022), but the time discretization is different. The simplified marker and cell method (Amsden and Harlow, 1970) is used to solve Eqs. (1) and (2). The implicit midpoint rule is used to discretize the time derivatives, and time marching is performed. The second-order central difference scheme is used for the discretization of the space derivatives. Similar to existing studies (Yanaoka and Inafune, 2023; Yanaoka, 2023), we used a simultaneous relaxation method for velocity and pressure. For the discretization of Eqs. (3) and (4), the implicit midpoint rule is applied for the time derivatives, and the fifth-order weighted essentially non-oscillatory (WENO) scheme (Jiang and Peng, 2000) is applied for the space derivatives. The existing research (Yanaoka and Nakayama, 2022) used the explicit method for the time derivatives in Eqs. (3) and (4), but considering the calculation efficiency, the time discretization was changed to the implicit method. In preparation for large-scale calculations to be performed in the future, we have improved the numerical method.
@@ 3.1 Instability analysis of a liquid ligament in a stationary fluid
As an initial condition in this calculation, similar to the previous study (Yanaoka and Nakayama, 2022), a disturbance with wavelength $\lambda$ and initial amplitude $\eta_{0}$ is given to the interface as follows:
@@
where $k=2\pi/\lambda$ is the wavenumber. The interface of the stationary ligament interface grows due to the influence of this disturbance. For the boundary conditions of the velocity and level set function, periodic boundary conditions are given at all boundaries.
@@
The liquid is a light oil and the surrounding gas is air. The densities of the liquid and gas are $\rho_{l}=851$ kg/m3 and $\rho_{g}=1.20$ kg/m3 and the viscosity coefficients are $\mu_{l}=2.225\times 10^{-3}$ Pa s and $\mu_{g}=1.827\times 10^{-5}$ Pa s, respectively. The surface tension coefficient is $\sigma=2.44\times 10^{-2}$ N/m. The reference values used in the calculation are $L_{\mathrm{ref}}=d$, $U_{\mathrm{ref}}=\sqrt{\sigma/(\rho_{l}d)}$, $\rho_{\mathrm{ref}}=\rho_{l}$, and $\mu_{\mathrm{ref}}=\mu_{l}$. The time intervals used in the calculation are $\Delta t/(d/U_{\mathrm{ref}})=1.0\times 10^{-3}$, $5.0\times 10^{-4}$, and $2.5\times 10^{-4}$ for grid1, grid2, and grid3, respectively. The time interval was set so that the Courant number was smaller than one.
@@ 3.2 Analysis of the deformation of liquid ligament interface in shear flow
Next, we investigate the effect of shear flow on the deformation of a liquid ligament. As shown in Fig. 1, the gas flows in the positive and negative directions of the $x$-axis at the velocity $U_{g}$ on the lower and upper sides of the liquid ligament, respectively. The velocity difference between the upper and lower boundaries is ${\Delta}U=2U_{g}$. In a liquid jet ejected from a nozzle, the direction of the shear velocity between gas and liquid is parallel to the central axis of a liquid column. On the other hand, in the model targeted by this study, the two directions are orthogonal, and therefore, we consider that the interface is likely to be deformed. We assumed a situation in which a ligament breaks up from the interface of a liquid jet or a liquid sheet, and the ligament is exposed to airflow. Such ligaments are present in non-uniform shear flows rather than uniform flows. Therefore, this study analyzes the behavior of a ligament in a shear flow.
As for the calculation region, as the liquid ligament is elongated in the $x$-direction due to the shear flow, the wide area in the $x$-direction is set. The region is $10d$, $8d$, and $\lambda$ in the $x$-, $y$-, and $z$-directions, respectively. The grid used for the calculations is the same as in Subsection 3.1.
@@
where $\delta$ is a parameter related to the velocity boundary layer thickness. In this study, we set $\delta/d=0.5$ and set the velocity boundary layer thickness to about the radius of the ligament so that the entire ligament is in an airflow with a velocity gradient. In addition, as the liquid flow in the high-speed flow field is turbulent, the formed liquid ligaments are not straight but curved (Marmottant and Villermaux, 2004b; Wahono et al., 2007). To model such a deformed liquid ligament, in this calculation, the position $Y_{c}$ of the central axis of the liquid ligament is changed in the $y$-direction as shown in the following equation:
@@
The computational domain is the same as the analysis in Subsection 3.2. We confirmed the grid dependence on calculation results using the same three grids as in Subsection 3.1. As a result, we found that valid results could be obtained with the grid resolution of grid2.
@@
where $\eta_{h}$, $\lambda_{h}$, and $k_{h}$ represent the amplitude, wavelength, and wavenumber of the high wavenumber component disturbance, respectively. The amplitude and wavenumber ratios of the high wavenumber component to the reference disturbance are defined as $\eta_{r}=\eta_{h}/\eta_{0}$ and $k_{r}=k_{h}/k$, respectively. When we give a single reference disturbance component to the liquid ligament interface, we set $\eta_{r}=0$ and $k_{r}=0$. Under this condition, we vary the wavenumber as $ka=$ 0.3, 0.5, 0.7, and 0.9. We investigated the effects of $\eta_{r}$ and $k_{r}$ on the deformation of the interface and found that $\eta_{r}=0.9$ had the greatest effect. In addition, when the wavenumbers of the high wavenumber disturbance were two and four times the wavenumbers of the low wavenumber disturbance, i.e., $k_{r}=$2.0 and 4.0, the effect of the high wavenumber disturbance appeared. In this calculation, we vary the wavenumber ratio as $k_{r}=$ 2.0 and 4.0 under $\eta_{r}=0.9$.
@@ 4.1 Instability analysis of liquid ligament in a stationary fluid
Figure 2(b) shows the dimensionless growth rate $\Omega$ in each grid compared with the theoretical value (Weber, 1931). At all wavenumbers, the calculated results agree well with the theoretical values.
@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
Next, we investigate the effect of shear flow on the deformation of the liquid ligament interface. Figure 3 shows the time variation of the flow field and liquid ligament for $\Delta U/U_{\mathrm{ref}}=11.9$ and $ka=0.7$. The velocity vector is the distribution in the $x$$y$ cross-section at $z/\lambda=0.01$. At $T=0-3$, the parts of $z/\lambda=0.5$ and $z/\lambda=$ 0 in the ligament are stretched by the shear flow in the positive and negative directions of the $x$-axis, respectively. Around the interface of the liquid ligament, the airflow flows along the interface, so the liquid ligament rotates counterclockwise with time. At $T=0.9$, we can see that the cross-sections at $z/\lambda=0$ and 0.5 in the liquid ligament become large.
@@
Figure 4 shows the interface and velocity vectors in the $x$$y$ cross-section at $T=0-0.3$. A solid black line indicates the interface. The left and right figures are the results at $z/\lambda=0.5$ and 1.0, respectively. At $T=0$, the liquid ligament is curved like a cosine function in the $z$-direction, and therefore, the lower interface ($y/d<0$) of $z/\lambda=0.5$ is exposed to faster airflow than the velocity near the upper interface ($y/d>0$). As a result, the liquid inside the liquid ligament is further accelerated, and at $T=0.1$, the liquid velocity around $(x/d,y/d)=(0.5,-0.5)$ is faster than that around $(x/d,y/d)=(-0.5,0.25)$. Similarly, in $z/\lambda=1.0$ at $T=0.1$, the liquid velocity around $(x/d,y/d)=(-0.5,0.5)$ is faster than that around $(x/d,y/d)=(0.5,-0.2)$. At the two cross-sections at $T=0.1$, vortices centered at $(x/d,y/d)=(0,0)$ are formed by the airflow. As the airflow flows along the interface, the vortex centers at the cross-sections of $z/\lambda=0.5$ and 1.0 move with time to the upper and lower sides of the interface, respectively. As a result, at $T=0.3$, the vortex centers exist near $(x/d,y/d)=(0.25,0.2)$ for $z/\lambda=0.5$ and near $(x/d,y/d)=(-0.25,-0.2)$ for $z/\lambda=1.0$.
@@
Figure 5 shows the interface, velocity vectors, and vorticity distribution in the $x$$z$ cross-section of $y/d=0$ and $y$$z$ cross-section of $x/d=0$ at $T=0.3$. The vorticity distributions in the $y$- and $x$-directions are shown in Figs. 5(a) and (b), respectively. In Fig. 4(d), the vortex centers in the cross-section of $z/\lambda=0.5$ and 1.0 exist at $y/d>0$ and $y/d<0$, respectively. Thus, at $z/\lambda=0.5$ and $z/\lambda=0$, 1.0 in Fig. 5(a), the airflows flow in the positive and negative directions of the $x$-axis at $y/d=0$. Around $(x/d,z/\lambda)=(1.0,0.2)$, $(-1.0,0.3)$ and $(x/d,z/\lambda)=(1.0,0.8)$, $(-1.0,0.7)$, counterclockwise and clockwise vortices are formed due to the difference in the direction of movement of the airflow, respectively. Similarly, around $(y/d,z/\lambda)=(0,0.25)$ and $(y/d,z/\lambda)=(0,0.75)$ in Fig. 5(b), clockwise and counterclockwise vortices are formed due to the deviation of the center positions of the vortices at $z/\lambda=0.5$ and 1.0 in Fig. 4(d). Thus, it is considered that the vortex caused by the airflow formed near the interface of the liquid ligament significantly influences the behaviors at the initial stage of the deformation of the liquid ligament.
@@
We investigate the changes in the flow field and liquid movement inside the liquid ligament due to the velocity difference at the time just before the liquid ligament breaks up. Figure 9 shows the interface, velocity vectors, vorticity contours, and streamlines at each $\Delta U/U_{\mathrm{ref}}$. Figure 9(a) is the result at $T=5.0$, and Figs. 9(b) and (c) are the results at $T=2.3$. In the left figure, the red solid line shows the streamline with a clockwise swirling component to the flow direction, and the blue solid line shows the reverse rotation. The velocity vector and $x$-direction vorticity distribution in Fig. 9(a) are the results of the $y$$z$ cross-section at $x/d=0$. The velocity vector and $y$-direction vorticity distribution in Figs. 9(b) and (c) are the results of the $x$$z$ cross-section at $y/d=0$. From the streamlines in Fig. 9(a), it can be seen that the liquid around $z/\lambda=0.5$ flows spirally toward $z/\lambda=0$ and 1.0. Due to such a three-dimensional flow, counter-rotating vortices are formed symmetrically to $y/d=0$ around $z/\lambda=0.3$ and 0.7 in the airflow. As these vortices induce the liquid movement along the central axis of the liquid ligament, the cross-sectional area of the liquid ligament decreases at $z/\lambda=0.5$ and increases at $z/\lambda=0$. Therefore, in Fig. 7, the droplet at $z/\lambda=0.5$ is smaller than that at $z/\lambda=0$.
In Fig. 9(b), the liquid existing near $z/\lambda=0.25$ and $z/\lambda=0.75$ flows toward $z/\lambda=0$, 0.5 and $z/\lambda=0.5$, 1.0, respectively. In addition, vortices near the interface develop. These vortices induce liquid movement along the central axis of the liquid ligament.

### v2

@@ Abstract
This study performed a numerical analysis of the deformation and breakup of a liquid ligament with various disturbances on the interface in shear flow. The shear flow generates a three-dimensional flow and vortices around the liquid ligament. These vortices promote the movement of the liquid inside the liquid ligament. When the velocity difference of shear flow increases, a nonlinear effect becomes strong, and turbulence with higher wavenumber components than the initial disturbance occurs at the interface. This turbulence accelerates the ligament splitting and increases the number of breakup droplets. Then, the droplet diameters become uniform, and the atomization quality improves. As the wavenumber of the disturbance applied to the interface increases, the liquid moving velocity along the central axis of the liquid ligament increases. Furthermore, the breakup time of the liquid ligament becomes short. In addition to the initial reference disturbance with a low wavenumber, when turbulence with twice the wavenumber of the reference disturbance is applied to the interface, the interface deformation and the splitting of the liquid ligament are similar to those with the single reference disturbance. When turbulence with four times the wavenumber is added to the initial disturbance with the low wavenumber at the interface, the deformation of the liquid ligament is accelerated, and the liquid ligament splits faster. Additionally, the number of breakup droplets increases; hence, the total surface area of the liquid increases. The droplet diameter becomes uniform; therefore, the atomization quality of the liquid ligament is improved.
@@ 1 INTRODUCTION
In the general atomization process, when the ejected liquid in the form of a liquid sheet or column is destabilized by the shearing of gas and liquid, a fine-scale liquid column, called a ligament, is generated, and the liquid eventually splits into small-scale droplets (Hashimoto, 1995). Numerous studies have been reported on the destabilization of liquid films, liquid columns, and droplets (Squire, 1953; O’Rourke and Amsden, 1987; Li and Tankin, 1991; Li et al., 2000; Kourmatzis and Masri, 2014; Fernandez et al., 2009; Shinjo and Umemura, 2010; Li et al., 2011; Odier et al., 2015; Gong et al., 2015). As spray consists of many small-scale droplets, it is considered that the behavior and splitting of liquid ligaments that generate droplets significantly affect the spray characteristics. However, the destabilization and breakup of liquid ligaments have not been elucidated in detail; therefore, further research focusing on the behavior of liquid ligaments is necessary.
@@
Marmottant and Villermaux (2004a) experimentally investigated the splitting behavior of liquid ligaments formed by pulling up a circular tube immersed in liquid and stretching the liquid. It has been clarified from this investigation that the ligament breakup time depends on the diameter of the tube, not on the speed with which the liquid is pulled up. Bizjan et al. (2014) conducted experiments on the breakup of liquid ligaments formed by centrifugal force using a rotating disk. Consequently, it has been clarified that the Weber number, Ohnesorge number, and liquid volumetric flow rate greatly influence the behavior and splitting length of the liquid ligament. In addition to the above studies, experimental (Henderson et al., 2000) and numerical (Dumouchel et al., 2015; Zhang and Yuan, 2022) studies on liquid ligaments have been reported. These studies did not focus on the deformation of liquid caused by gas–liquid shear velocity and did not investigate the effect of surrounding airflows on the behavior of liquid ligaments. Constante-Amores et al. (2021) captured a ligament around a liquid jet interface with direct numerical simulation. The ligament detached from the interface and formed droplets. In this manner, fine-scale liquids such as ligaments split in atomization processes. Therefore, the breakup characteristics of ligaments should be investigated to clarify spray characteristics. The authors (Yanaoka and Nakayama, 2022) performed a three-dimensional analysis of the destabilization and splitting of a liquid ligament in shear flow. When the wavenumber of initial disturbance increases, the breakup time of the ligament in the shear flow becomes short, and the diameter of the formed droplet becomes small. Furthermore, as the velocity difference of the shear flow increases, the dispersion of the droplet diameter decreases, and the droplet size becomes uniform.
@@
Airflows exist around liquids injected from injectors. As liquid ligaments are destabilized by gas–liquid shear velocity, it is necessary to investigate the effect of shear flow on the deformation and splitting of liquid ligaments. Additionally, both air and liquid flows are turbulent in high-speed jets. In such a turbulent flow field, the velocity turbulence intensity increases around the gas–liquid interface. At this time, disturbances with various wavenumber components occur at the interface, and it is considered that these turbulences promote the deformation and splitting of liquid ligaments. However, it is difficult to experimentally investigate the effects of various turbulences on interface deformation, focusing on fine-scale liquid ligaments generated in a turbulent flow. Hence, the influences of wide-ranging wavenumber disturbances on interface deformation have not been fully elucidated. Thus, it is necessary to investigate the effects of various disturbances on the behavior of liquid ligaments.
@@ 2 NUMERICAL PROCEDURES
This study considers incompressible viscous flow. We apply the level set method (Sussman et al., 1994) to track the interface. Additionally, the continuum surface force method (Brackbill et al., 1992) is used to evaluate surface tension, which is treated as a body force. The governing equations for incompressible two-phase flow are the continuity equation, the Navier-Stokes equation, the advection equation for the level set function, and the reinitialization equation, which are given as follows:
@@
where $t$ is the time, ${\bf u}=(u,v,w)$ is the velocity vector at the coordinates ${\bf x}=(x,y,z)$, $\rho$ is the density, $p$ is the pressure, $\mu$ is the viscosity coefficient, and ${\bf D}$ is the strain rate tensor. $\kappa$ is the interface curvature, $\hat{\bf n}$ is the unit normal vector at the interface, $\delta_{s}$ is the delta function, and $\phi$ is the level set function. $\tau$ is a pseudo time for reinitialization, and $\phi_{0}$ is the level set function before reinitialization. The variables in the fundamental equations are nondimensionalized as follows using the reference values of length $L_{\mathrm{ref}}$, velocity $U_{\mathrm{ref}}$, density $\rho_{\mathrm{ref}}$, and viscosity coefficient $\mu_{\mathrm{ref}}$:
@@
The numerical method is almost the same as the previous research (Yanaoka and Nakayama, 2022), but the time discretization is different. The simplified marker and cell method (Amsden and Harlow, 1970) is used to solve Eqs. (1) and (2). The implicit midpoint rule is used to discretize the time derivatives, and time marching is performed. The second-order central difference scheme is used for the discretization of the space derivatives. Similarly to existing studies (Yanaoka and Inafune, 2023; Yanaoka, 2023), we used a simultaneous relaxation method for velocity and pressure. For the discretization of Eqs. (3) and (4), the implicit midpoint rule is applied for the time derivatives, and the fifth-order weighted essentially nonoscillatory (WENO) scheme (Jiang and Peng, 2000) is applied for the space derivatives. The existing research (Yanaoka and Nakayama, 2022) used the explicit method for the time derivatives in Eqs. (3) and (4), but considering the calculation efficiency, the time discretization was changed to the implicit method. In preparation for large-scale calculations to be performed in the future, we have improved the numerical method.
@@ 3.1 Instability analysis of a liquid ligament in a stationary fluid
As an initial condition in this calculation, similarly to the previous study (Yanaoka and Nakayama, 2022), a disturbance with wavelength $\lambda$ and initial amplitude $\eta_{0}$ is given to the interface as follows:
@@
where $k=2\pi/\lambda$ is the wavenumber. The interface of the stationary ligament interface grows owing to the influence of this disturbance. For the boundary conditions of the velocity and level set function, periodic boundary conditions are given at all boundaries.
@@
The liquid is a light oil and the surrounding gas is air. The densities of the liquid and gas are $\rho_{l}=851$ kg/m3 and $\rho_{g}=1.20$ kg/m3 and the viscosity coefficients are $\mu_{l}=2.225\times 10^{-3}$ Pa s and $\mu_{g}=1.827\times 10^{-5}$ Pa s, respectively. The surface tension coefficient is $\sigma=2.44\times 10^{-2}$ N/m. The reference values used in the calculation are $L_{\mathrm{ref}}=d$, $U_{\mathrm{ref}}=\sqrt{\sigma/(\rho_{l}d)}$, $\rho_{\mathrm{ref}}=\rho_{l}$, and $\mu_{\mathrm{ref}}=\mu_{l}$. The time intervals used in the calculation are $\Delta t/(d/U_{\mathrm{ref}})=1.0\times 10^{-3}$, $5.0\times 10^{-4}$, and $2.5\times 10^{-4}$ for grid1, grid2, and grid3, respectively. The time interval is set such that the Courant number is smaller than one.
@@ 3.2 Analysis of the deformation of liquid ligament interface in shear flow
Subsequently, we investigate the effect of shear flow on the deformation of a liquid ligament. As shown in Fig. 1, the gas flows in the positive and negative directions of the $x$-axis at the velocity $U_{g}$ on the lower and upper sides of the liquid ligament, respectively. The velocity difference between the upper and lower boundaries is ${\Delta}U=2U_{g}$. In a liquid jet ejected from a nozzle, the direction of the shear velocity between gas and liquid is parallel to the central axis of a liquid column. Conversely, in the model targeted by this study, the two directions are orthogonal; thus, it is considered that the interface is likely to be deformed. We assumed a situation in which a ligament breaks up from the interface of a liquid jet or a liquid sheet, and the ligament is exposed to airflow. Such ligaments are present in nonuniform shear flows rather than uniform flows. Hence, this study analyzes the behavior of a ligament in a shear flow.
As for the calculation region, as the liquid ligament is elongated in the $x$-direction owing to the shear flow, the wide area in the $x$-direction is set. The region is $10d$, $8d$, and $\lambda$ in the $x$-, $y$-, and $z$-directions, respectively. The grid used for the calculations is the same as in Subsection 3.1.
@@
where $\delta$ is a parameter related to the velocity boundary layer thickness. In this study, we set $\delta/d=0.5$ and set the velocity boundary layer thickness to about the radius of the ligament so that the entire ligament is in an airflow with a velocity gradient. Additionally, as the liquid flow in the high-speed flow field is turbulent, the formed liquid ligaments are not straight but curved (Marmottant and Villermaux, 2004b; Wahono et al., 2007). To model such a deformed liquid ligament, in this calculation, the position $Y_{c}$ of the central axis of the liquid ligament is changed in the $y$-direction as shown in the following equation:
@@
The computational domain is the same as the analysis in Subsection 3.2. We confirmed the grid dependence on calculation results using the same three grids as in Subsection 3.1. Consequently, we found that valid results could be obtained with the grid resolution of grid2.
@@
where $\eta_{h}$, $\lambda_{h}$, and $k_{h}$ represent the amplitude, wavelength, and wavenumber of the high wavenumber component disturbance, respectively. The amplitude and wavenumber ratios of the high wavenumber component to the reference disturbance are defined as $\eta_{r}=\eta_{h}/\eta_{0}$ and $k_{r}=k_{h}/k$, respectively. When we give a single reference disturbance component to the liquid ligament interface, we set $\eta_{r}=0$ and $k_{r}=0$. Under this condition, we vary the wavenumber as $ka=$ 0.3, 0.5, 0.7, and 0.9. We investigated the effects of $\eta_{r}$ and $k_{r}$ on the deformation of the interface and found that $\eta_{r}=0.9$ had the greatest effect. Additionally, when the wavenumbers of the high wavenumber disturbance were two and four times the wavenumbers of the low wavenumber disturbance, i.e., $k_{r}=$2.0 and 4.0, the effect of the high wavenumber disturbance appeared. In this calculation, we vary the wavenumber ratio as $k_{r}=$ 2.0 and 4.0 under $\eta_{r}=0.9$.
@@ 4.1 Instability analysis of liquid ligament in a stationary fluid
Figure 2(b) shows the dimensionless growth rate $\Omega$ in each grid compared with the theoretical value (Weber, 1931). At all wavenumbers, the calculated results support the theoretical values.
@@ 4.2 Analysis of the deformation of liquid ligament interface in shear flow
Subsequently, we investigate the effect of shear flow on the deformation of the liquid ligament interface. Figure 3 shows the time variation of the flow field and liquid ligament for $\Delta U/U_{\mathrm{ref}}=11.9$ and $ka=0.7$. The velocity vector is the distribution in the $x$$y$ cross-section at $z/\lambda=0.01$. At $T=0-3$, the parts of $z/\lambda=0.5$ and $z/\lambda=$ 0 in the ligament are stretched by the shear flow in the positive and negative directions of the $x$-axis, respectively. Around the interface of the liquid ligament, the airflow flows along the interface, so the liquid ligament rotates counterclockwise with time. At $T=0.9$, we can see that the cross-sections at $z/\lambda=0$ and 0.5 in the liquid ligament become large.
@@
Figure 4 shows the interface and velocity vectors in the $x$$y$ cross-section at $T=0-0.3$. A solid black line indicates the interface. The left and right figures are the results at $z/\lambda=0.5$ and 1.0, respectively. At $T=0$, the liquid ligament is curved like a cosine function in the $z$-direction; thus, the lower interface ($y/d<0$) of $z/\lambda=0.5$ is exposed to faster airflow than the velocity near the upper interface ($y/d>0$). Consequently, the liquid inside the liquid ligament is further accelerated, and at $T=0.1$, the liquid velocity around $(x/d,y/d)=(0.5,-0.5)$ is faster than that around $(x/d,y/d)=(-0.5,0.25)$. Similarly, in $z/\lambda=1.0$ at $T=0.1$, the liquid velocity around $(x/d,y/d)=(-0.5,0.5)$ is faster than that around $(x/d,y/d)=(0.5,-0.2)$. At the two cross-sections at $T=0.1$, vortices centered at $(x/d,y/d)=(0,0)$ are formed by the airflow. The airflow flows along the interface; thus, the vortex centers at the cross-sections of $z/\lambda=0.5$ and 1.0 move with time to the upper and lower sides of the interface, respectively. Consequently, at $T=0.3$, the vortex centers exist near $(x/d,y/d)=(0.25,0.2)$ for $z/\lambda=0.5$ and near $(x/d,y/d)=(-0.25,-0.2)$ for $z/\lambda=1.0$.
@@
Figure 5 shows the interface, velocity vectors, and vorticity distribution in the $x$$z$ and $y$$z$ cross-sections of $y/d=0$ and $x/d=0$, respectively, at $T=0.3$. The vorticity distributions in the $y$- and $x$-directions are shown in Figs. 5(a) and (b), respectively. In Fig. 4(d), the vortex centers in the cross-section of $z/\lambda=0.5$ and 1.0 exist at $y/d>0$ and $y/d<0$, respectively. Thus, at $z/\lambda=0.5$ and $z/\lambda=0$, 1.0 in Fig. 5(a), the airflows flow in the positive and negative directions of the $x$-axis at $y/d=0$. Around $(x/d,z/\lambda)=(1.0,0.2)$, $(-1.0,0.3)$ and $(x/d,z/\lambda)=(1.0,0.8)$, $(-1.0,0.7)$, counterclockwise and clockwise vortices are formed by the difference in the direction of movement of the airflow, respectively. Similarly, around $(y/d,z/\lambda)=(0,0.25)$ and $(y/d,z/\lambda)=(0,0.75)$ in Fig. 5(b), clockwise and counterclockwise vortices are formed by the deviation of the center positions of the vortices at $z/\lambda=0.5$ and 1.0 in Fig. 4(d). Therefore, it is considered that the vortex caused by the airflow near the interface of the liquid ligament significantly influences the behaviors at the initial stage of the deformation of the liquid ligament.
@@
We investigate the changes in the flow field and liquid movement inside the liquid ligament owing to the velocity difference at the time just before the liquid ligament breaks up. Figure 9 shows the interface, velocity vectors, vorticity contours, and streamlines at each $\Delta U/U_{\mathrm{ref}}$. Figure 9(a) is the result at $T=5.0$, and Figs. 9(b) and (c) are the results at $T=2.3$. In the left figure, the red solid line shows the streamline with a clockwise swirling component to the flow direction, and the blue solid line shows the reverse rotation. The velocity vector and $x$-direction vorticity distribution in Fig. 9(a) are the results of the $y$$z$ cross-section at $x/d=0$. The velocity vector and $y$-direction vorticity distribution in Figs. 9(b) and (c) are the results of the $x$$z$ cross-section at $y/d=0$. From the streamlines in Fig. 9(a), it can be seen that the liquid around $z/\lambda=0.5$ flows spirally toward $z/\lambda=0$ and 1.0. Because of such a three-dimensional flow, counter-rotating vortices are formed symmetrically to $y/d=0$ around $z/\lambda=0.3$ and 0.7 in the airflow. As these vortices induce the liquid movement along the central axis of the liquid ligament, the cross-sectional area of the liquid ligament decreases at $z/\lambda=0.5$ and increases at $z/\lambda=0$. Therefore, in Fig. 7, the droplet at $z/\lambda=0.5$ is smaller than that at $z/\lambda=0$.
In Fig. 9(b), the liquid existing near $z/\lambda=0.25$ and $z/\lambda=0.75$ flows toward $z/\lambda=0$, 0.5 and $z/\lambda=0.5$, 1.0, respectively. Additionally, vortices near the interface develop. These vortices induce liquid movement along the central axis of the liquid ligament.